not sure how to prove this?

if n is odd then n^2 = 1 (mod 4)

thanks!

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3838)


n = an odd integer
n = 2k+1 for some integer k
n^2 = (2k+1)^2
n^2 = 4k^2+4k+1 after using the FOIL rule
n^2 = 4*(k^2+k) + 1
n^2 = 4*integer + 1
The last equation shows that whatever n^2 happens to be, it's 1 more than a multiple of 4. This only applies when n is odd.

We have shown that (n^2)/4 gives some quotient remainder 1 when n is odd.
And also proves n^2 = 1 (mod 4) when n is odd.

Some examples.
nn^2(n^2)/4n^2 (mod 4)
110 remainder 11
392 remainder 11
5256 remainder 11
74912 remainder 11
98120 remainder 11
1112130 remainder 11

Try some other examples out for yourself.
When doing modular arithmetic, we ignore the quotient to focus on the remainder only.

Extra Credit Question: Try to prove that n^2 = 0 (mod 4) when n is even. A hint is that n = 2k.

Answer by ikleyn(53937)

.
not sure how to prove this?

if n is odd then n^2 = 1 (mod 4)

thanks!
~~~~~~~~~~~~~~~~~~~~


                It is easy.


If n is odd integer number, it can be presented in the form  n = 2k+1,  
where k is some integer number.


Then 

    n^2 = (2k)^2 + 2*(2k) + 1 = 4k^2 + 4k + 1.    (1)


In the right side, first addend, 4k^2, is a multiple of 4, 
and the second addend 4k is a multiple of k.


Hence, the sum of the first two addends is a multiple of k.


Thus n^2 = 1 (mod 4).

At this point, the proof is complete.


In whole, this all is at the level of self-evident truths.